SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)
A Brief History of EMC

Please feel free to add stuff and/or remove stale items. This file is
$BFROOT/www/Detector/Calorimeter/Operations/pages/EMChistory.html.
Also, don't forget to log any changes to at least the EMC section of the logbook.

  Source Calibration

  • look at: /BFROOT/www/Detector/Calorimeter/RadSource/index.html

  •   Other Calibrations

  • June 5, 2002: (Gary&Ivo) Electronics Calib for Crate 3 - no constants loaded, only for diagnosing changes in source cal constants.

  • March 29, 2002: (Gary) Electronics Calib for Crate 0

  • March 26, 2002: (Gary) Pedestal Calib for Crate 0

  • February 26, 2002: Bhabha Calib

  •   Electronics Changes

  • June 15, 2002: (Ingrid) Replaced LV PS in BF06 (Noisy channels developed ... move when swapping LVPS to new slot).
  • June 13, 2002: (Ingrid) 8 LV PS spares put in BF00-BF07 for testing (12 hour MD). No problems found, so they are left there. Old good ones now in spare cabinet.
  • May 22, 2002: (Ingrid) LV PS (BF04) still had problems. Found that the new PS put in yesterday was also bad. Replaced with another (good!) spare.

  • May 21, 2002: (Ingrid) LV PS (BF04) Swithced off ... power-cycled, now ok.

  • May 8, 2002 (Ingrid) TRB FB2 and FB3 have been swapped for diagnostics for the DAQ-crate3, slot0 problem.
  • April 26, 2002 (Jong) LV PS - swapped BB06 and BB07 (diagnostics)

  • April 09, 2002: (Ingrid) LV PS (BF04) Replacement

  • March 26, 2002: (Ingrid) LV PS (BF09) Replacement

  • March 26, 2002: (Ingrid) LV PS (BF01 and BF02) Swapped (for diagnostics... and left there)

  • March 26, 2002: (Ingrid/Gary/Jong) Bias for Minicrate 01 in BF Q0 placed on channel 10 and B-diodes masked out

  •   Other Changes/Implementation

  • July 3, 2002: (Jong) Temprature sensor cables from board 5, Quad2, which have been swapped in the early days, this has now been corrected in software (Epics).
  • June 10, 2002: (Ingrid) HVAC fixes the EC fluorinert chiller. Spare chiller has a bad leak. This is being fixed too, by cleaning a seal.
  • June 8, 2002: (Jong) EC fluorinert chiller stopped. Compressor-fan pump broken. Spare chiller used instead.
  • May 2, 2002 (David, Eric & Ingrid) Replaced the RTD downstream of heater in water chiller supply.
  • May 2, 2002: (Jong)
  • Bias Voltage ramping under injection has been ready. We're not running with it yet.
  • April 12, 2002: (Gary) Switched to range change correction LUT as of run 27584.

  • April 9, 2002: (Ingrid)
  • EMC PIN diode offset adjusted so that EMC_TOTL reading went from 0.2 to 0.11.
  • March 26, 2002 (Gerard et. al.) New water pump installed for water chiller.
  • March 22, 2002: (Jong)
  • Pedestals for leakage currents implemented. Bias/Background/OdcBackground epics panels have been updated.


    This file is $BFROOT/www/Detector/Calorimeter/Operations/pages/EMChistory.html
    Page author: Jong Yi
    Last Modified: Kitty Cats of SLAC
    Last modified: Tue Jun 18 16:32:28 PDT 2002